OpenStudy (anonymous):

8/11 what is it in decimal form

OpenStudy (anonymous):

0.72 repeating

OpenStudy (unklerhaukus):

do you know how to use a calculator/

OpenStudy (anonymous):

8/11 = 72/99 = 0.727272...

OpenStudy (anonymous):

i do but i want to talk to people online like you

OpenStudy (anonymous):

money bird it is only 72 repeating